# pnasresearcharticle

### v1

@@ Abstract
peakedness |||| smart card data |||| human mobility habit |||| departure time
2]4
@@
Human mobility habits are a reflection of how our daily lives are framed by a series of regularised activities. These activities are each located in space and time and act to constrain how, where, when and why we move around. When considered in aggregate across a transport system exert important implications for congestion (1), environmental pollution (2) and the travel experience (3).
@@
Scholarship points to the way that daily travel patterns are repetitive, and individuals tend to recurrently follow the same travel choices (4). We know that individual trip making is conditioned by a range of factors such as gender (5), age (6), occupation (7), household characteristics (8), socio-economics (9) and is further governed by characteristics of the built environment (10). However much of this work tends to be cross sectional in nature and does not attempt to capture the longer-term histories of individual travellers that is an essential ingredient to understand travel habits (4). Thus, if we are to first capture and then empirically measure the way in which travel habits impact our transit systems from which smart solutions to invoke ‘good’ and ‘sustainable’ travel habits can be implemented then we must begin to assemble a longitudinal dimension to travel behaviour scholarship.
@@
Smart card systems employed to automate fare collection have emerged as valuable source of information on travel behaviour with important strategic, tactical, and operational benefits (11). Studies have exploited these data to investigate travel demand (12), trip purpose (13), geographic patterns (14), temporal travel patterns (15), community structure (16), service reliability (17), modal transfer behaviour (18), and the passenger experience (19), to name a select few. Critically, as it relates to understating travel habits, these data offer a continuous record of transit usage through which we can move beyond the traditional cross-sectional studies and use of travel diaries to assemble longitudinal histories of individual travellers.
@@
In this study we draw on a large smart card database of bus ridership and extend the work on travel behaviour by examining long term travel trajectories for individual passengers across a metropolis. More specifically, we develop a novel measurement framework to unveil travel habits by examining the departure time choice behaviours of individual bus riders over a 12-month period. In the transport literature, travel habits has been largely studied in the context of travel mode choice (20, 21, 22, 23) and route choice (24, 25, 26, 27), but research in the context of departure time choice (28) is rare. We aim to fill this gap by developing a new approach to measure habitual aspects of day-to-day departure time distribution in terms of bus riders’ tendency towards choosing the same departure time repeatedly—that we term ‘peakedness’.
@@
Departure time distribution, which shows the number of trip departures across different times of the day, provides valuable insights into when and how many trips are initiated. This information is crucial for network operators, transport planners, and service providers to understand peak demand periods, enabling them to align resources, services, and infrastructure to meet varying travel demands throughout the day. The departure time distribution can be constructed for an individual traveller by combining day-to-day departure time choices over a certain period (e.g., 1 year) or can be constructed at the system-wide by combining departure time distributions from all the travellers in the system, forming the system-wide demand loading profile, like city-wide departure time distributions shown in Fig.1A. While traditional statistics like standard deviation or kurtosis can provide insights into the shape of a distribution by measuring the ‘spread’ or ‘tailedness’ characteristics, they do not directly measure the ‘peakedness’ of the distribution (see Supporting Information (SI) Section 1.A). In our study, we introduce indicators specifically designed to capture the ‘peakedness’ characteristics of the departure time distribution. Moreover, we derive an analytical expression that offer an intuitive explanation on how the ‘peakedness’ of individual travellers’ departure time distributions affect the ‘peakedness’ of the overall system-wide departure time distribution. These indicators are aimed at providing a more robust and systematic approach to analysing the degree of concentration of trips during specific times of the day. The concept of ‘peakedness’ and associated measurement framework offer a new and important lens through which to examine travel habits of individual users across a metropolis and how these habits scale to form metropilitan-wide dynamics.
@@
Figure 1: Bus passenger trips in three major cities in Queensland, Australia, namely Brisbane, Gold Coast, and Sunshine Coast. [A] city-wide departure time distributions for bus commuters, constructed from a 12-month (July 2015 to June 2016) record of go card, an electronic smart card ticketing system; [B, C, D] locations of bus stops
Results
A Measure of Departure Time Peakedness
@@
Eq.[1] finds the time window $[t,t+h]$ that contains the largest proportion of departure times, and the corresponding interval probability $\psi_{h}$ represents the proportion of trips that fall within the most peaked $h$-minute window. Fig.2 illustrates this concept, where departure time distributions are shown in the period of 6-8AM, peaked around 7AM. Then, the area under the density curve within the window of $h$-minutes centring at 7AM represents the value of the $h$-min Peak Trip Concentration, $\psi_{h}$. When a departure time distribution is highly peaked, this proportion is larger and, thus, the peakedness of the distribution is characterised as a larger $\psi_{h}$ value (Fig.2A). When a departure time distribution is less peaked, its peakedness level is lower and characterised as a smaller $\psi_{h}$ value (Fig.2B). The value of $\psi_{h}$ ranges between 0 and 1, where $\psi_{h}$ is 1 (the largest) when all departure times happen within the $h$-min peak window and $\psi_{h}$ is close to 0—more accurately as low as $\psi_{h}=h/1440$ (e.g., $\psi_{h}=15/1440=0.01$ if $h$ is set to 15 minutes)—when the departure times are uniformly distributed over the whole day (24 hours = 1440 minutes).
Although the proposed metric fundamentally measures the probability of events occurring within a defined window—something not new in probability theory—, the novelty lies in its application as a peakedness metric for departure time distributions. Traditional measures such as standard deviation, kurtosis, or entropy assess the overall characteristics of a distribution but may not fully capture how sharply or consistently departure times are concentrated within a specific interval (see SI Section 1.A). By quantifying the probability that departures cluster around a specific peak time, the proposed metric offers a robust measure of behaviour that is less sensitive to outliers. It is well-supported by both theory and practice that measuring the concentration of activity within the most intense time period provides a meaningful way to quantify the peakedness of a distribution. Examples that share conceptual similarities with the proposed metric are relative peakedness in probability theory and Peak Hour Factor (PHF) in traffic engineering (see SI Sections 1.B-1.D. Additionally, using a probability-based measure allows us to link individual peakedness metrics to system-wide peakedness. Since the system-wide departure time distribution can be expressed as a mixture of individual distributions, the proposed metric enables the exploration of how individual behavioural patterns shape the overall system’s departure time patterns. Therefore, while the concept of measuring probabilities is well established, its application to assess individual peakedness in departure time distributions represents a novel and practical contribution.
@@
Figure 2: Measuring the $h$-min Peak Trip Concentration, $\psi_{h}$, as an indicator of the peakedness of a departure time distribution. The area under the density curve within the $h$-min peak window represents $\psi_{h}$: [A] the value of $\psi_{h}$ is larger when the distribution is more peaked and [B] the value of $\psi_{h}$ is smaller when the distribution is less peaked.
@@
A Measurement Framework for the Peakedness of Travellers’ Departure Time Choices and Their Relations to System-wide Dynamics
@@
Fig.3 provides a visual representation of the relationship among these three measures in three different scenarios. In Fig.3A-B, we illustrate a situation where the system-wide departure time distribution exhibits high peakedness (high $\psi_{h}^{sys}$) as a result of both the individual users’ departure time distributions being highly peaked (high $\overline{\psi_{h}}$) and the alignment of their peak periods being high (high $PCF_{h}$). Conversely, the other two cases in the figure showcase scenarios where the system-wide departure time has low peakedness. Fig.3C-D illustrates a situation where this low peakedness arises from a lack of coincidence in the users’ peak periods (low $PCF_{h}$), despite each user individually having a highly peaked departure time distribution (high $\overline{\psi_{h}}$). Fig.3E-F, on the other hand, illustrates a situation where the low system-wide peakedness results from the individual users’ departure time distributions having low peakedness (low $\overline{\psi_{h}}$), despite their peak periods being highly aligned (high $PCF_{h}$).
@@
Measuring Departure Time Peakedness of Three Cities

### v2

# Understanding everyday public transit travel habits: a measurement framework for the peakedness of departure time distributions

@@ Abstract
peakedness , smart card data , travel habit , departure time , public transport , human mobility
journal: Travel Behaviour and Society
@@
[inst1]organization=School of Civil Engineering, The University of Queensland,addressline=St Lucia, city=Brisbane, postcode=4072, state=Queensland, country=Australia
[…]
[inst2]organization=School of the Environment, The University of Queensland,addressline=St Lucia, city=Brisbane, postcode=4072, state=Queensland, country=Australia
1 Introduction
@@
Human mobility habits are a reflection of how our daily lives are framed by a series of regularised activities. These activities are each located in space and time and act to constrain how, where, when and why we move around. When considered in aggregate across a transport system exert important implications for congestion (Mannering et al., 1994), environmental pollution (Schwanen et al., 2012) and the travel experience (Domarchi et al., 2008).
@@
Scholarship points to the way that daily travel patterns are repetitive, and individuals tend to recurrently follow the same travel choices (Gärling and Axhausen, 2003). We know that individual trip making is conditioned by a range of factors such as gender (Gordon et al., 1989), age (Chudyk et al., 2015), occupation (Rasouli et al., 2015), household characteristics (Dieleman et al., 2002), socio-economics (Kotval-K and Vojnovic, 2015) and is further governed by characteristics of the built environment (Hong et al., 2014). However much of this work tends to be cross sectional in nature and does not attempt to capture the longer-term histories of individual travellers that is an essential ingredient to understand travel habits (Gärling and Axhausen, 2003). Thus, if we are to first capture and then empirically measure the way in which travel habits impact our transit systems from which smart solutions to invoke ‘good’ and ‘sustainable’ travel habits can be implemented then we must begin to assemble a longitudinal dimension to travel behaviour scholarship.
@@
Smart card systems employed to automate fare collection have emerged as valuable source of information on travel behaviour with important strategic, tactical, and operational benefits (Pelletier et al., 2011). Studies have exploited these data to investigate travel demand (Morency et al., 2007), trip purpose (Lee and Hickman, 2014), geographic patterns (Tao et al., 2014), temporal travel patterns (Cats and Ferranti, 2022), community structure (Yildirimoglu and Kim, 2018), service reliability (Ma et al., 2015), modal transfer behaviour (Sun et al., 2015), and the passenger experience (Chu and Lomone, 2016), to name a select few. Critically, as it relates to understating travel habits, these data offer a continuous record of transit usage through which we can move beyond the traditional cross-sectional studies and use of travel diaries to assemble longitudinal histories of individual travellers.
@@
In this study we draw on a large smart card database of bus ridership and extend the work on travel behaviour by examining long term travel trajectories for individual passengers across a metropolis. More specifically, we develop a novel measurement framework to unveil travel habits by examining the departure time choice behaviours of individual bus riders over a 12-month period. In the transport literature, travel habits has been largely studied in the context of travel mode choice (Bamberg et al., 2003; Gardner, 2009; Cherchi and Manca, 2011; Sharmeen and Timmermans, 2014) and route choice (Bogers et al., 2005; Kurauchi et al., 2014; He et al., 2014; Kim et al., 2017), but research in the context of departure time choice (Thorhauge et al., 2020) is rare. We aim to fill this gap by developing a new approach to measure habitual aspects of day-to-day departure time distribution in terms of bus riders’ tendency towards choosing the same departure time repeatedly—that we term ‘peakedness’. This ‘peakedness’ shares conceptual similarities with the ‘stickiness’ of route choice behaviour defined by Kim et al. (2017), which captures the tendency of bus riders to consistently choose or ‘stick to’ the same route. While the ‘stickiness index’ introduced by Kim et al. (2017) could theoretically be applied to departure time choice, it is primarily designed for ‘discrete’ route choice alternatives, motivating us to devise a new measure that directly addresses the ‘continuous’ nature of departure time choice.
@@
Departure time distribution, which shows the number of trip departures across different times of the day, provides valuable insights into when and how many trips are initiated. This information is crucial for network operators, transport planners, and service providers to understand peak demand periods, enabling them to align resources, services, and infrastructure to meet varying travel demands throughout the day. The departure time distribution can be constructed for an individual traveller by combining day-to-day departure time choices over a certain period (e.g., 1 year) or can be constructed at the system-wide by combining departure time distributions from all the travellers in the system, forming the system-wide demand loading profile. Figure 1 shows examples of city-wide departure time distributions for bus commuters, constructed using a 12-month smart card data from three major cities in Queensland, Australia, namely Brisbane, Gold Coast, and Sunshine Coast. While traditional statistics like standard deviation or kurtosis can provide insights into the shape of a distribution by measuring the ‘spread’ or ‘tailedness’ characteristics, they do not directly measure the ‘peakedness’ of the distribution. In our study, we introduce indicators specifically designed to capture the ‘peakedness’ characteristics of the departure time distribution. Moreover, we derive an analytical expression that offer an intuitive explanation on how the ‘peakedness’ of individual travellers’ departure time distributions affect the ‘peakedness’ of the overall system-wide departure time distribution. These indicators are simple and easy to apply, yet insightful, aiming to provide a more robust and systematic approach to analysing the degree of concentration of trips during specific times of the day. The concept of ‘peakedness’ and associated measurement framework offer a new and important lens through which to examine travel habits of individual users across a metropolis and how these habits scale to form metropolitan-wide dynamics.
@@
Figure 1: Bus passenger trips in three major cities in Queensland, Australia, namely Brisbane, Gold Coast, and Sunshine Coast: city-wide departure time distributions for bus commuters are constructed from a 12-month (July 2015 to June 2016) record of go card, an electronic smart card ticketing system.
2 Methodology
2.1 A Measure of Departure Time Peakedness
@@
Eq.[1] finds the time window $[t,t+h]$ that contains the largest proportion of departure times, and the corresponding interval probability $\psi_{h}$ represents the proportion of trips that fall within the most peaked $h$-minute window. Figure 2 illustrates this concept, where departure time distributions are shown in the period of 6-8AM, peaked around 7AM. Then, the area under the density curve within the window of $h$-minutes centring at 7AM represents the value of the $h$-min Peak Trip Concentration, $\psi_{h}$. When a departure time distribution is highly peaked, this proportion is larger and, thus, the peakedness of the distribution is characterised as a larger $\psi_{h}$ value (Figure 2(a)). When a departure time distribution is less peaked, its peakedness level is lower and characterised as a smaller $\psi_{h}$ value (Figure 2(b)). The value of $\psi_{h}$ ranges between 0 and 1, where $\psi_{h}$ is 1 (the largest) when all departure times happen within the $h$-min peak window and $\psi_{h}$ is close to 0—more accurately as low as $\psi_{h}=h/1440$ (e.g., $\psi_{h}=15/1440=0.01$ if $h$ is set to 15 minutes)—when the departure times are uniformly distributed over the whole day (24 hours = 1440 minutes).
@@
Figure 2: Measuring the $h$-min Peak Trip Concentration, $\psi_{h}$, as an indicator of the peakedness of a departure time distribution. The area under the density curve within the $h$-min peak window represents $\psi_{h}$: (a) the value of $\psi_{h}$ is larger when the distribution is more peaked and (b) the value of $\psi_{h}$ is smaller when the distribution is less peaked.
To measure departure time peakedness at the individual level, the $h$-min Peak Trip Concentration, $\psi_{h,i}$, is estimated for each user $i$. Specifically, $\psi_{h,i}$ is calculated for each user for a specific origin-destination (OD) pair to evaluate each traveller’s own departure time consistency within their individual day-to-day trip distributions for recurring trips. By analysing departure time peakedness for each unique user-OD combination separately, we effectively control for variations in travel distances across users and ODs, ensuring that our method captures how consistently an individual sticks to their preferred departure time window.
Although the proposed metric fundamentally measures the probability of events occurring within a defined window—something not new in probability theory—, the novelty lies in its application as a peakedness metric for departure time distributions. Traditional measures such as standard deviation, kurtosis, or entropy assess the overall characteristics of a distribution but may not fully capture how sharply or consistently departure times are concentrated within a specific interval (see A). By quantifying the probability that departures cluster around a specific peak time, the proposed metric offers a robust measure of behaviour that is less sensitive to outliers. It is well-supported by both theory and practice that measuring the concentration of activity within the most intense time period provides a meaningful way to quantify the peakedness of a distribution. Examples that share conceptual similarities with the proposed metric are relative peakedness in probability theory (see B) and Peak Hour Factor (PHF) in traffic engineering (see C). Additionally, using a probability-based measure allows us to link individual peakedness metrics to system-wide peakedness. Since the system-wide departure time distribution can be expressed as a mixture of individual distributions, the proposed metric enables the exploration of how individual behavioural patterns shape the overall system’s departure time patterns. Therefore, while the concept of measuring probabilities is well established, its application to assess individual peakedness in departure time distributions represents a novel and practical contribution.
In this study, we focus on departure time distributions rather than arrival time distributions, as departure times more directly reflect individual behavioural preferences and habitual travel choices, whereas arrival times are influenced by additional external factors such as congestion and travel time variability. By capturing the repeated selection of preferred departure time windows within each traveller’ day-to-day trip choices, our approach provides a clearer measure of how consistently individuals adhere to their habitual travel schedules. Furthermore, a comparative analysis of departure and arrival time peakedness across users, presented in E, shows a strong correlation between the two measurement approaches, demonstrating that despite variations in travel time, the overall peakedness profile remains consistent between departure and arrival time distributions.
@@
2.2 A Measurement Framework for the Peakedness of Travellers’ Departure Time Choices and Their Relations to System-wide Dynamics
@@
Figure 3 provides a visual representation of the relationship among these three measures in three different scenarios. In Figure 3(a)-(b), we illustrate a situation where the system-wide departure time distribution exhibits high peakedness (high $\psi_{h}^{sys}$) as a result of both the individual users’ departure time distributions being highly peaked (high $\overline{\psi_{h}}$) and the alignment of their peak periods being high (high $PCF_{h}$). Conversely, the other two cases in the figure showcase scenarios where the system-wide departure time has low peakedness. Figure 3(c)-(d) illustrates a situation where this low peakedness arises from a lack of coincidence in the users’ peak periods (low $PCF_{h}$), despite each user individually having a highly peaked departure time distribution (high $\overline{\psi_{h}}$). Figure 3(e)-(f), on the other hand, illustrates a situation where the low system-wide peakedness results from the individual users’ departure time distributions having low peakedness (low $\overline{\psi_{h}}$), despite their peak periods being highly aligned (high $PCF_{h}$).
@@
2.3 Choice of Peak Window, $h$
Selecting the appropriate width of the peak window, $h$, is important in measuring and interpreting $\psi_{h}$, especially when comparing its values across individual users. If $h$ is too large, $\psi_{h}$ will tend to approach 1, as most departures fall within the wide window, leading to an inflated peakedness for most travellers. Conversely, if $h$ is too small, $\psi_{h}$ will approach 0, since very few departures times fall within the narrow window, resulting in artificially low peakedness for most travellers. A desirable property of $h$ is that the resulting $\psi_{h}$ effectively differentiate between different users’ varying departure time patterns. As such, we propose to determine the optimal window $h^{*}$ as the $h$ value that maximises the variance of $\psi_{h}$ ($\text{Var}[\psi_{h}]$) across users as follows:
where $\psi_{h,i}$ is the $h$-min peak trip concentration of user $i$, $\overline{\psi_{h}}$ is the sample mean of $\psi_{h}$ over $n$ users, and $\text{Var}[\psi_{h}]$ is the variance of $\psi_{h}$, which can be estimated ($\mathrel{\hat{=}}$) by calculating the sample variance. Figure 4(a) illustrates how the variance of $\psi_{h}$ changes with respect to different choice of peak window $h$. When $h$ is very small, the variance is low because $\psi_{h}\approx 0$ for most users. As $h$ increases, the variance initially rises but eventually decreases again when $h$ becomes too large and $\psi_{h}\approx 1$ for most users. Therefore, the variance of $\psi_{h}$ as a function of $h$ forms a peak at a critical point, $h=h^{*}$, after which it decreases. This critical point, $h=h^{*}$, serves as the optimal window width that maximises the spread of $\psi_{h}$ values across users, thereby best distinguishing different levels of peakedness in the system.
In a special case where a system has only two peakedness levels—one group of users with highly peaked departure time distribution $f_{1}$ and another group of users with less peaked departure time distribution $f_{2}$—, the variance-maximising optimal window $h^{*}$ corresponds to the peak window at which the two distribution curves intersect, as indicated by the two vertical dotted lines in Figure 4(b). The detailed proof is provided in D. This example offers a useful intuition that the variance-maximising optimal $h^{*}$ captures the best peak window that naturally separate different peakedness levels as much as possible, which offers a good middle ground for $h$ that is not too small and not too large given the set of departure time distributions in the dataset.
Figure 4: Optimal peak window: (a) the change in the variance of $\psi_{h}$ across users with respect to different choice of $h$; (b) an illustration of the variance-maximising optimal peak window $h^{*}$ in a special case where a system has two peakedness levels, i.e., departure time distribution with high peakedness ($f_{1}$) and departure time distribution with low peakedness ($f_{2}$)
3 Results
3.1 Data
The three peakedness measures developed in this work—$\psi_{h}$, $\psi_{h}^{sys}$, and $PCF_{h}$—are estimated using public transport smart card data from South East Queensland (SEQ), Australia, to characterise the departure time peakedness of bus passengers. This study considers three different study networks, namely, Brisbane, Gold Coast, and Sunshine Coast, which are three major cities in the SEQ region, as shown in Figure 5.
(a) Brisbane
(b) Gold Coast
(c) Sunshine Coast
Figure 5: Maps of Brisbane, Gold Coast, and Sunshine Coast showing the locations of bus stops (blue dots)
We use a 12-month (July 2015 to June 2016) record of go card, an electronic smart card ticketing system used on the public transport network in SEQ operated by TransLink (the public transit agency for Queensland). The go card is the principal means by which passengers pay their transit fares in SEQ, representing approximately 87% of all trips taken across the TransLink public transport network (TransLink, 2016). It requires passengers to touch the card on a card reader at the start and end of each journey as well as when transferring between trips, where TransLink defines a trip as the act of travelling from point A to point B with no transfers (a single trip) and a journey as the act of travelling from the origin to the final destination that may involve one or a number of trips (transfers) using different transport modes (TransLink, 2021). The go card data contain information on each trip’s boarding and alighting locations and their associated times, trip ID, journey ID, card ID (go card identifier), and passenger type, among other attributes. Passenger types recorded in the go card database include Adult, Senior/Pensioner, Tertiary Student, Secondary Student (secondary school students aged 15+ years), Child (children aged 5 to 14 years), and others (special concession card holders such as veterans).
In this study, we focus on regular bus passengers who use buses repeatedly. For each bus user, trip data are processed and grouped to form journeys for each origin-destination (OD). Then, the departure time distribution is constructed for each user-OD pair by extracting the boarding times at the origin stop for a given OD journey over the course of one year from July 2015 to June 2016. The user-OD pair is, thus, the basic unit for measuring the peakedness of user-level departure time distribution, i.e., $h$-min peak trip concentration for user $i$ ($\psi_{h,i}$). Each departure time distribution is represented as a normalised histogram with a bin width of 5 minutes, as shown in the examples in Figure 6.
Figure 6 shows examples of three bus users’ departure time distributions and the estimated $\psi_{h}$ for a peak window width of $h=15$ minutes. The $\psi_{h}$ values for Users 1, 2, and 3 are 0.843, 0.445, and 0.152, respectively. User 1 has the strongest tendency towards choosing a particular departure time as the great majority of his/her trips (84.3%) occur during the short 15-min peak window. User 2 shows more flexibility in choosing departure time, resulting in about half his/her trips (44.5%) occurring during the 15-min peak. User 3 has the lowest peakedness and the greatest flexibility among three, showing that his/her most peaked 15-minutes only covers 15.2% of the trips and the remaining 84.4% of data are widely spread out from this peak.
Table 1 shows the summary of dataset size for the three study networks after filtering in terms of the number of users, user-OD pairs, journeys, and origins, with and without grouping by passenger type. The entire data encompasses 5,947,907 bus journeys made by 29,640 individuals across the three study networks. ‘The number of users’ in the table represents the total count of all distinct go card users in each dataset. Since each user may have multiple ODs and, thus, result in multiple user-OD pairs, the number of distinct user-OD pairs is greater than the number of users. ‘The number of journeys’ is the sum of the number of journeys across all user-OD pairs. ‘The number of origins’ is the number of unique bus stops from which these journeys started. Overall, the Brisbane dataset is the largest, which has 50,336 user-OD pairs providing 50,336 departure time distributions and their associated ($\psi_{h}$) measures, followed by Gold Coast and Sunshine Coast, which has 3968 and 1817 user-OD pairs, respectively. In terms of passenger type, Adult is the largest group in all three cities. The second largest group differs, however, as Tertiary Student is the second largest group in Brisbane and Gold Coast, whereas Senior/Pensioner is the second largest group in Sunshine Coast.
Table 1: Summary of datasets
3.2 Measuring Departure Time Peakedness of Three Cities
In this section, we measure and compare the departure time peakedness of the three study networks. Figure 1 shows the system-wide departure time distribution for each network, combining all bus commuters’ journeys over a one-year period. In Brisbane, two distinct peaks (AM and PM) are observed. Gold Coast and Sunshine Coast display similar dual-peak patterns, although the PM peaks are less sharp compared to Brisbane. Given these two distinct peaks, we separately analyse the peakedness characteristics of the AM and PM periods.
For each user-OD pair, we first construct trip frequency distributions using 5-minute bins by aggregating trip observations over multiple days (1 year in this study). We then identify the peak 5-minute bin (i.e., the time interval with the highest trip density) for each user-OD pair. The user-OD departure time distributions with their peak 5-minute bin occurring between 12:00 AM and 11:59 AM are classified as AM trips, while those with their peak occurring between 12:00 PM and 11:59 PM are classified as PM trips. For each time-of-day group, we estimate three peakedness measures—$\overline{\psi_{h}}$, $\psi_{h}^{sys}$, and $PCF_{h}$—using the associated user-OD departure time distributions. Individual $h$-min peak trip concentration ($\psi_{h,i}$) is first computed for each user-OD pair by summing trip densities over an $h$-minute interval centred around the peak 5-minute bin. To obtain the average user-level peakedness ($\overline{\psi_{h}}$), we then take the mean of $\psi_{h,i}$ across all user-OD pairs within a given network and time-of-day group. Separately, we aggregate all individual user-OD departure time data to construct the system-level mixture distribution and estimate system-level peakedness ($\psi_{h}^{sys}$) by summing densities over an $h$-minute interval centred around the peak 5-minute bin of this distribution. Finally, $PCF_{h}$ is derived using Eq. [6].